If α is a repeated root of ax2 + bx + c = 0, then Limx→α tan(ax2 + bx + c) / (x - α)2 is ???

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear Mainak

α is a repeated root of ax2 + bx + c = 0

so   ax2 + bx + c=a(x-α)2

now  Limx→α tan(ax2 + bx + c) / (x - α)

         =Limx→α [tana(x-α)2 ]/ (x - α)2

        =Limx→α a[tana(x-α)2 ]/ a(x - α)2

        =a             because   Limx→0 tanx /x   = 1 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin